# Gauss Principle in Incompressible Flow: A Unified Variational Perspective on Pressure and Projection

### v1

@@ Abstract
Following recent work (Gonzalez & Taha 2022; Peters & Ormiston 2025), this manuscript clarifies what the Gauss–Appell principle determines in incompressible, inviscid flow and how it connects to classical projection methods. At a fixed time, freezing the velocity and varying only the material acceleration leads to minimization of a quadratic subject to acceleration-level constraints. First-order conditions yield a Poisson–Neumann problem for a reaction pressure whose gradient removes the non-solenoidal and wall-normal content of the provisional residual, i.e. the Leray–Hodge projection. Thus, Gauss–Appell enforces the instantaneous kinematic constraints and recovers Euler at the instant. In steady flows, this principle cannot select circulation or stagnation points because these are properties of the velocity state, not the instantaneous acceleration correction. The principle only determines the reaction pressure for an already-specified velocity field. The impressed/reaction pressure decomposition can be supplemented with orthogonality conditions, allowing physical inputs (e.g. circulation, freestream) to be embedded while maintaining exact constraint enforcement. This variational viewpoint provides a diagnostic for computational incompatibility: In a practical computation, spikes in projection effort may signal problematic boundary conditions or under-resolution, and clarifies how pressure instantaneously maintains solenoidality without creating or destroying vorticity. The goal of this note is simply to lend more clarity to the application of the Gauss principle, and to connect it concretely to well known concepts including potential flow theory, projection algorithms, and recent variational approaches.
@@
The application of variational principles to fluid mechanics has a rich history spanning over 150 years, beginning with Helmholtz’s vortex theorems (1858) and Kelvin’s circulation theorem (1869). Classical approaches typically employed Hamilton’s principle, extremizing action integrals over time to derive the equations of motion through calculus of variations (Holm et al., 1998; Truesdell & Noll, 2004; Marchioro & Pulvirenti, 2012). These time-integral formulations, while mathematically elegant, vary entire space-time paths and thus provide limited insight into the instantaneous mechanics of constraint enforcement in incompressible flows.  (Holm et al., 1998; Truesdell & Noll, 2004; Marchioro & Pulvirenti, 2012).
A fundamentally different variational approach emerged from Gauss’s principle of least constraint (Gauss, 1829), originally developed for constrained mechanical systems. Unlike Hamilton’s principle which varies paths over time, Gauss’s principle operates at a fixed instant, minimizing a quadratic form in accelerations subject to acceleration-level constraints. This instantaneous character makes it particularly suited to understanding how pressure maintains incompressibility moment by moment. Despite its potential, the principle remained largely unexplored in fluid mechanics for over a century. Recent works (Gonzalez & Taha, 2022; Taha & Shorbagy, 2025; Peters & Ormiston, 2025) have brought back a variational lens on fluid flows, particularly revisiting the Gauss principle in the context of the airfoil lift problem.
@@
Gonzalez & Taha (2022) minimize a steady reduced functional on a kinematically admissible manifold (e.g., a potential‑flow family parameterized by circulation), a legitimate model‑reduction device but not the instantaneous Gauss principle. Taha & Shorbagy (2025) recast this as a “minimum pressure‑gradient” criterion as a selection mechanism for weak Euler solutions—again a steady or quasi‑steady choice layered on top of the equations. Peters & Ormiston (2025) correctly argue that least constraint, by itself, cannot resolve classical indeterminacies of steady exterior potential flow.
This work reconciles these viewpoints by showing precisely where Gauss/Appell ends (fixed‑time projection) and where additional modeling/selection begins (steady‑state choice of manifold or boundary data). The Gauss principle is a fixed-time variational statement: one freezes the velocity field and varies the material acceleration. Minimizing a positive‑definite “Appellian” in the admissible accelerations yields first order optimality conditions that are exactly the Leray–Hodge projection of the residual onto gradient fields; the Lagrange multiplier is a reaction pressure obtained from a Poisson–Neumann problem whose gradient removes the non‑solenoidal part of the residual. In steady flow, this principle does not select among kinematically admissible steady states; it only determines the pressure compatible with the given velocity field. These points are well known from projection methods but are often obscured in recent discussions; we make the equivalence and its scope precise.
Further, the non-uniqueness of the impressed vs. constraint split is clarified, and the decomposition of Peters & Ormiston (2025) is substantiated via an energy-orthogonality condition that selects a unique decomposition relative to a chosen impressed subspace, and by connecting this result to classical projection. It is shown however, that while physically intuitive choices can be made for the impressed subspace, the choice itself is arbitrary.
Unlike the principle of least action, Gauss/Appell is a least‑deviation principle in the accelerations at a given instant. In this sense, it is clarified that Gauss/Appell provides a unifying variational interpretation of the pressure Poisson solve—useful analytically and practically, particularly when one wishes to embed physical inputs (e.g., circulation or far‑field content) through an impressed subspace while keeping the constraints exact. We present rigorous derivations in both Eulerian and Lagrangian, and provide insight into the physical implications of the Gauss principle.
@@ 1 Setting, notations
Let $\Omega$ be the exterior domain of a smooth body, and $\partial\Omega$ be the boundaries. The density $\rho$ is constant. $u(\cdot,t)$ is the velocity field. The Gauss’ principle is applied at a fixed time. One freezes the state $u(\cdot,t)$ and chooses an admissible instantaneous update in acceleration. For fluids, the material acceleration is $a\triangleq\partial_{t}u+(u\!\cdot\!\nabla)u.$ At a fixed time, the convective part $(u\!\cdot\!\nabla)u$ is a known function of space (because $u(\cdot,t)$ is frozen). Hence a perturbation of acceleration comes entirely from perturbing $u_{t}$: $\delta a=\delta u_{t}$. This is the fluid analogue of Gauss in constrained multibody dynamics (Udwadia & Kalaba, 1992), where one minimizes a positive-definite quadratic form in accelerations at a fixed configuration and velocity, subject to acceleration-level constraints; here the Eulerian fixed-time data are $(u,\nabla\!\cdot u,u\!\cdot n)$. Therefore, the variational variable is $u_{t}(\cdot,t)$, not a time-slab history, and not a configuration displacement as in stationary action.
@@
Below, the Eulerian form of the derivation is detailed. The Lagrangian form can be found in the appendix.
@@ 1.1 Impressed vs reaction pressure and the Appellian
As in Peters & Ormiston (2025), we split the total pressure
@@
  • $P_{F}$ (impressed) collects the gradient content we treat as “given” at the instant and that does not necessarily enforce the kinematic constraints.
@@ 2 First order stationarity
Construct the Lagrangian with multipliers $\lambda$ (volume) and $\mu$ (boundary):
@@
together with (1). This derivation is aligned with  Peters & Ormiston (2025), though we clarify the frozen field C.
@@
This is precisely the standard projection method logic in continuous form: solve a Poisson–Neumann problem for pressure and project the residual onto gradients. It is notable that the Leray projector (see, for instance, Temam (2024)) maps any field to the divergence-free, wall-tangent subspace. Writing $R\triangleq\rho\ a+\nabla P_{F}=\rho(u_{t}+C)+\nabla P_{F},$ the KKT conditions $\rho(u_{t}^{\star}+C)=-\nabla P_{R}$ says $R=-\nabla P_{R}$: the solenoidal, wall-tangent part of $R$ vanishes. Thus $P_{R}$ realizes the Leray projection of $R$ onto gradients, and 5 is the strong-form projector.
@@ 3 Physical implications and insight
At a fixed instant, the Gauss–Appell step has a concrete physical meaning: it chooses the smallest admissible change to the material acceleration that restores the instantaneous momentum balance while preserving incompressibility and wall non-penetration. The only agency it has is a reaction field, $\nabla P_{R}/\rho$, whose role is to remove whatever divergence and normal flux the provisional residual carries. The pressure is a reaction that instantaneously enforces the kinematics; it does no work on divergence-free, wall-tangent motions. This point is correctly argued by  Gonzalez & Taha (2022). After elimination, the minimized Appellian $\mathcal{A}^{\ast}$ becomes a quantitative measure of the projection effort, i.e. how hard the constraint must “push” to keep the instantaneous update compatible with incompressibility and wall tangency.
@@ 3.1 Steady flow
seeking $\Gamma^{\star}=\arg\min_{\Gamma}S(\Gamma)$. This is a legitimate reduced-model optimization on a one-parameter manifold of kinematically admissible steady flows and can, in practice, pick a $\Gamma$ consistent with certain steady configurations. It can be used as a closure condition. However, it is not the instantaneous Gauss/Appell principle; it presupposes steadiness and a particular parameterization. In the full Gauss setting, the state $u$ is frozen, the variable is $u_{t}$, and the outcome is the constraint projection $P_{R}$2), not the selection of $\Gamma$. In the steady case, for every $\Gamma$ there exists a pressure that makes $(u(\cdot;\Gamma),p(\cdot;\Gamma))$ a steady Euler solution. Applying Gauss at the instant then yields $u_{t}^{\ast}=0$ for every $\Gamma$. Therefore, the principle does not (and cannot) decide which $\Gamma$ nature takes. In fact, in the steady flow case, the application of the Gauss principle does not yield any new information other than the fact that the reaction pressure can be determined.
Taha & Shorbagy (2025) propose the ‘Principle of Minimum pressure gradient’, and offer useful reduced‑manifold selections that align with classical expectations (Kutta condition on nearly sharp airfoils, Glauert’s rapidly rotating cylinder, empirical separation angles). An important distinction is required to interpret scope: Gauss/Appell provides the fixed‑time projection; selection over steady states requires additional modeling choices.
In the unsteady case, since $u_{t}$ (hence $a$) is varied with $u$ held fixed, acceleration-level constraints are enforced, resulting in the reaction pressure via the Poisson–Neumann problem (5). This exactly recovers Euler at the instant and clarifies the roles of impressed vs constraint pressure. As argued by  Peters & Ormiston (2025), this also explains why least constraint cannot, by itself, resolve the classical indeterminacy of steady exterior potential flow. That indeterminacy lives in the state manifold, not in the instantaneous constrained acceleration update.
3.2 Impressed/constraint split is arbitrary
Let $\psi$ be a smooth field (e.g. $\psi=\Gamma\,\psi_{s}$ representing a circulation mode). Define a new split $\widetilde{P}_{F}\triangleq P_{F}+\psi,\quad\widetilde{P}_{R}\triangleq P_{R}-\psi.$ Then
@@ 3.2 Impressed/constraint split is arbitrary
Similarly, the Poisson equation changes to $\nabla^{2}\widetilde{P}_{R}=-\rho\,\nabla\!\cdot C-\nabla^{2}\psi,\ \partial_{n}\widetilde{P}_{R}=-\rho\,C\!\cdot n-\partial_{n}\psi$. Therefore $\widetilde{P}_{R}=P_{R}-\psi+\text{const}$, exactly as above. So the Poisson step leaves $(u_{t}^{\star},p)$ invariant. Thus, in incompressible Euler at a fixed time, any decomposition $p=P_{F}+P_{R}$ leaves momentum balance unchanged. This simple observation allows one to assign an impressed component $P_{F}$ and let the constraint $P_{R}$ enforce the acceleration‑level constraints, with no change to the physical solution. Once $P_{R}$ is solved, the optimal update is $u_{t}^{\star}=-(C+\nabla P_{R}/\rho)$, and the instantaneous Euler balance $\rho(\partial_{t}u+(u\!\cdot\!\nabla)u)=-\nabla(P_{F}+P_{R})$ is recovered.

### v2

@@ Abstract
Following recent work (Gonzalez & Taha, 2022; Peters & Ormiston, 2025), this manuscript clarifies what the Gauss–Appell principle determines in incompressible, inviscid flow and how it connects to classical projection methods. At a fixed time, freezing the velocity and varying only the material acceleration leads to minimization of a quadratic subject to acceleration-level constraints. First-order conditions yield a Poisson–Neumann problem for a reaction pressure whose gradient removes the non-solenoidal and wall-normal content of the provisional residual, i.e. the Leray–Hodge projection. Thus, Gauss–Appell enforces the instantaneous kinematic constraints and recovers Euler at the instant. In steady flows, this principle –on its own– cannot select circulation or stagnation points because these are properties of the velocity state, not the instantaneous acceleration correction. The principle only determines the reaction pressure for an already-specified velocity field. The impressed/reaction pressure bookkeeping can be supplemented with orthogonality conventions that separate prescribed conservative forcing (if any) from the reaction enforcing the constraints. This variational viewpoint also yields a simple computational diagnostic: the minimized Appellian equals a $L^{2}$ norm of the reaction-pressure gradient which vanishes for constraint-compatible updates and grows with the magnitude of divergence and wall-flux mismatch. The goal of this note is simply to lend more clarity to the application of the Gauss principle, and to connect it concretely to well known concepts including potential flow theory, recent variational approaches and projection algorithms.
@@
The application of variational principles to fluid mechanics has a rich history spanning over 150 years, beginning with Helmholtz’s vortex theorems (1858) and Kelvin’s circulation theorem (1869). Classical approaches typically employed Hamilton’s principle, extremizing action integrals over time to derive the equations of motion through calculus of variations (Holm et al., 1998; Truesdell & Noll, 2004; Marchioro & Pulvirenti, 2012). These time-integral formulations, while mathematically elegant, vary entire space-time paths and thus provide limited insight into the instantaneous mechanics of constraint enforcement in incompressible flows  (Holm et al., 1998; Truesdell & Noll, 2004; Marchioro & Pulvirenti, 2012).
A fundamentally different variational approach emerged from Gauss’s principle of least constraint (Gauss, 1829), originally developed for constrained mechanical systems. Unlike Hamilton’s principle which varies paths over time, Gauss’s principle operates at a fixed instant, minimizing a quadratic form in accelerations subject to acceleration-level constraints. This instantaneous character makes it particularly suited to understanding how pressure maintains incompressibility moment by moment. Despite its potential, the principle remained relatively uncommon in classical fluid mechanics for over a century, even though it is widely used in constrained molecular dynamics (e.g. Nosé (1984); Hoover (1985); Evans & Morris (1990); Thijssen (2007). Recent works (Gonzalez & Taha, 2022; Taha et al., 2023; Taha & Shorbagy, 2025; Peters & Ormiston, 2025) have brought back a variational lens on fluid flows, particularly revisiting the Gauss principle in the context of the airfoil lift problem.
@@
Gonzalez & Taha (2022) minimize a steady reduced functional on a kinematically admissible manifold (e.g., a potential‑flow family parameterized by circulation), a model‑reduction approach but not the instantaneous Gauss principle. Taha et al. (2023) derive a general “minimum pressure-gradient” (PMPG) interpretation directly from Gauss’s principle (for arbitrary viscous stresses), showing that Navier–Stokes arises as the first-order optimality condition when the variation is taken with respect to the local acceleration. Steady/inviscid specializations of such costs have also been explored as selection mechanisms layered on top of Euler (e.g. Taha & Shorbagy (2025)). In particular, the fixed-time Gauss/Appell projection step considered here does not, by itself, resolve classical indeterminacies of steady exterior potential flow, because those indeterminacies live in the choice of steady velocity state rather than in the instantaneous constrained acceleration correction (see also Peters & Ormiston (2025)).
This work clarifies these viewpoints by showing precisely where Gauss/Appell ends (fixed‑time projection) and where additional modeling/selection begins (steady‑state choice of manifold or boundary data). The Gauss principle is a fixed-time variational statement: one freezes the velocity field and varies the material acceleration. Minimizing a positive‑definite “Appellian” in the admissible accelerations yields first order optimality conditions that are exactly the Leray–Hodge projection of the residual onto gradient fields; the Lagrange multiplier is a reaction pressure obtained from a Poisson–Neumann problem whose gradient removes the non‑solenoidal part of the residual. In steady flow, this principle does not select among kinematically admissible steady states; it only determines the pressure compatible with the given velocity field. These points are well known from projection methods but are often obscured in recent discussions; we make the equivalence and its scope precise.
Further, we clarify that shifting a known scalar potential between the two terms in $p=P_{F}+P_{R}$ is a gauge freedom in bookkeeping and does not imply physical non-uniqueness of the reaction (constraint) pressure. Building on Peters & Ormiston (2025), we show how an energy (Dirichlet) orthogonality condition can be used as a gauge-fixing convention to pick a unique representative once a reference/impressed subspace is chosen, and connect this to the classical Leray projection. When $P_{F}$ is reserved for truly impressed physics (e.g. conservative body-force potentials specified independently of incompressibility), the split is unique and $P_{R}$ is the unique Lagrange multiplier enforcing the constraints (up to a constant).
Unlike the principle of least action, Gauss/Appell is a least‑deviation principle in the accelerations at a given instant. In this sense, it is clarified that Gauss/Appell provides a unifying variational interpretation of the pressure Poisson solve—useful analytically and practically, particularly when one wishes to embed physical inputs (e.g., circulation or far‑field content) through an impressed subspace while keeping the constraints exact. We present rigorous derivations and provide insight into the physical implications of the Gauss principle.
@@ 1 Setting, notations
Let $\Omega$ be the exterior domain of a smooth body, and $\partial\Omega$ be the boundaries. The density $\rho$ is constant. $u(\cdot,t)$ is the velocity field. The Gauss’ principle is applied at a fixed time. One freezes the state $u(\cdot,t)$ and chooses an admissible instantaneous update in acceleration. For fluids, the material acceleration is $a\triangleq\partial_{t}u+(u\!\cdot\!\nabla)u.$ We assume throughout that $u(\cdot,t)$ is sufficiently smooth such that the quadratic functional below is finite. At a fixed time, the convective part $(u\!\cdot\!\nabla)u$ is a known function of space (because $u(\cdot,t)$ is frozen). Hence a perturbation of acceleration comes entirely from perturbing $u_{t}$: $\delta a=\delta u_{t}$. This is the fluid analogue of Gauss in constrained multibody dynamics (Udwadia & Kalaba, 1992), where one minimizes a positive-definite quadratic form in accelerations at a fixed configuration and velocity, subject to acceleration-level constraints; here the Eulerian fixed-time data are $(u,\nabla\!\cdot u,u\!\cdot n)$. Therefore, the variational variable is $u_{t}(\cdot,t)$, not a time-slab history, and not a configuration displacement as in stationary action.
@@
Below, the Eulerian form of the derivation is detailed.
@@ 1.1 Impressed vs reaction pressure and the Appellian
Motivated by Peters & Ormiston (2025), we write the total pressure
@@
  • $P_{F}$ (impressed / prescribed) is a scalar potential associated with conservative impressed forcing that is specified independently of the incompressibility and wall constraints (e.g. a body-force potential); in unforced Euler one may take $P_{F}\equiv 0$.
@@
For a fixed velocity snapshot $u(\cdot,t)$ and prescribed impressed potential $P_{F}$, the reaction pressure $P_{R}$ is determined uniquely (up to an additive constant) by the Poisson–Neumann problem (5). In the linearized unforced setting emphasized in Peters & Ormiston (2025), taking the divergence yields $\nabla^{2}p=0$, so only the sum $p=P_{F}+P_{R}$ is constrained to be harmonic. The stronger separation $\nabla^{2}P_{F}=\nabla^{2}P_{R}=0$ (their Eq. (17)) is therefore an additional harmonic-gauge convention rather than a restriction on physically impressed forces, which are prescribed independently of the incompressibility constraint. In the present work we do not infer such restrictions on $P_{F}$; once any impressed conservative forcing is specified, the reaction pressure $P_{R}$ enforcing (1) is uniquely determined (up to a constant) by (5).
@@
The quadratic form in (3) uses the mass-weighted $L^{2}$ inner product on accelerations, and yields the classical Leray–Hodge projection. In the function-space setting this choice is not unique: other metrics define different least-constraint projections. Weaker norms and lower regularity settings have been used in variational approaches to high-Reynolds-number flows; see e.g. Hoffman & Johnson (2010); Hoffman et al. (2016).
@@ 2 First order stationarity
We write the constrained Lagrangian and derive the first order Karush-Kuhn-Tucker (KKT) conditions. The Lagrangian with multipliers $\lambda$ (volume) and $\mu$ (boundary) is:
@@
together with (1). This derivation follows the same constrained-optimization/projection-step logic as Peters & Ormiston (2025), while emphasizing that in the present work $P_{F}$ denotes only prescribed conservative impressed forcing (if any) and the reaction pressure $P_{R}$ carries all constraint-enforcing pressure content.
@@
This reduced functional is the instantaneous $L^{2}$ norm of the reaction pressure gradient (the “minimum pressure-gradient” quantity emphasized by Taha et al. (2023)). This is precisely the standard projection method logic in continuous form: solve a Poisson–Neumann problem for pressure and project the residual onto gradients. It is notable that the Leray projector (see, for instance, Temam (2024)) maps any field to the divergence-free, wall-tangent subspace.
Writing $R\triangleq\rho\ a+\nabla P_{F}=\rho(u_{t}+C),$ the KKT conditions $\rho(u_{t}^{\star}+C)=-\nabla P_{R}$ says $R=-\nabla P_{R}$: the solenoidal, wall-tangent part of $R$ vanishes. Thus $P_{R}$ realizes the Leray projection of $R$ onto gradients, and 5 is the strong-form projector.
@@ 3 Physical implications and insight
At a fixed instant, the Gauss–Appell step has a concrete physical meaning: it chooses the smallest admissible change to the material acceleration that restores the instantaneous momentum balance while preserving incompressibility and wall non-penetration. The only agency it has is a reaction field, $\nabla P_{R}/\rho$, whose role is to remove whatever divergence and normal flux the provisional residual carries. The pressure is a reaction that instantaneously enforces the kinematics; it does no work on divergence-free, wall-tangent motions. This point is correctly argued by  Gonzalez & Taha (2022).
3.1 Projection effort as a diagnostic
After elimination, the minimized Appellian $\mathcal{A}^{\ast}$ becomes a quantitative measure of the projection effort, i.e. how hard the constraint must “push” to keep the instantaneous update compatible with incompressibility and wall tangency. The minimized Appellian can be written as
$\mathcal{A}_{\star}=\frac{1}{2\rho}\,\|\nabla P_{R}\|_{L^{2}(\Omega)}^{2},$
and therefore provides a scalar measure of the magnitude of the constraint correction at a given instant. In particular, if the frozen residual $C$ is already compatible with the acceleration constraints, $\nabla\!\cdot C=0$ in $\Omega$ and $C\!\cdot n=0$ on $\partial\Omega$, then (5) reduces to Laplace’s equation with homogeneous Neumann data and hence $\nabla P_{R}\equiv 0$ (up to an additive constant), so that $\mathcal{A}_{\star}=0$. In computations, a sudden growth of $\mathcal{A}_{\star}$ indicates that the provisional update violates divergence or wall-flux constraints and the projection is applying a large correction; this may occur, for example, due to inconsistent boundary treatment or under-resolution. More details on the projection operator in the context of general boundary conditions can be found in Maria Denaro (2003).
@@ 3.2 Steady flow
seeking $\Gamma^{\star}=\arg\min_{\Gamma}S(\Gamma)$. This is a reduced-order optimization on a one-parameter manifold of kinematically admissible steady flows and appears to pick a $\Gamma$ consistent with the Kutta condition in nearly sharp airfoils. However, it does not directly follow from the instantaneous Gauss/Appell principle; it presupposes steadiness and a particular parameterization. Peters & Ormiston (2025) discuss cases where such criteria can select steady states that do not align with physical reality.
In the full Gauss setting, the state $u$ is frozen, the variable is $u_{t}$, and the outcome is the constraint projection $P_{R}$2), not the selection of $\Gamma$. In the steady case, for every $\Gamma$ there exists a pressure that makes $(u(\cdot;\Gamma),p(\cdot;\Gamma))$ a steady Euler solution. Applying Gauss at the instant then yields $u_{t}^{\ast}=0$ for every $\Gamma$. Therefore, the direct fixed-time application of Gauss’s principle does not, by itself, favor a particular value of $\Gamma$; it simply returns the reaction pressure associated with each already-specified steady velocity state. In the steady case, the fixed-time Gauss step adds no additional closure beyond computing the compatible pressure (any selection over steady states requires additional modeling/criteria).
Taha et al. (2023) (see also Taha & Shorbagy (2025)) propose the ‘Principle of Minimum pressure gradient’, and offer reduced‑manifold selections that align with classical expectations from applying the Kutta condition on nearly sharp airfoils. An important distinction is required to interpret scope: Gauss/Appell provides the fixed‑time projection; selection over steady states requires additional modeling choices.
In the unsteady case, since $u_{t}$ (hence $a$) is varied with $u$ held fixed, acceleration-level constraints are enforced, resulting in the reaction pressure via the Poisson–Neumann problem (5). This exactly recovers Euler at the instant and clarifies the roles of impressed vs constraint pressure. This also explains why the fixed-time least-constraint step does not, by itself, resolve the classical indeterminacy of steady exterior potential flow: the indeterminacy lives in the state manifold, not in the instantaneous constrained acceleration update (cf. discussion in Peters & Ormiston (2025)). That indeterminacy lives in the state manifold, not in the instantaneous constrained acceleration update.
3.3 Freedom in the impressed/reaction pressure bookkeeping
Let $\psi$ be any prescribed smooth scalar field (a reference pressure / gauge potential). Define a new split $\widetilde{P}_{F}\triangleq P_{F}+\psi,\quad\widetilde{P}_{R}\triangleq P_{R}-\psi.$ Then
@@ 3.3 Freedom in the impressed/reaction pressure bookkeeping
Similarly, the Poisson equation changes to $\nabla^{2}\widetilde{P}_{R}=-\rho\,\nabla\!\cdot C-\nabla^{2}\psi,\ \partial_{n}\widetilde{P}_{R}=-\rho\,C\!\cdot n-\partial_{n}\psi$. Therefore $\widetilde{P}_{R}=P_{R}-\psi+\text{const}$, exactly as above. So the Poisson step leaves $(u_{t}^{\star},p)$ invariant. This invariance reflects the classical fact that adding a prescribed gradient force (or, equivalently, redefining a reference pressure) does not affect the divergence-free, wall-tangent evolution; it is a gauge freedom in the representation of $p$, not a physical non-uniqueness of the reaction force. Once $P_{F}$ is fixed by the impressed physics (if any), the reaction pressure $P_{R}$ solving (5) is uniquely determined up to an additive constant. Once $P_{R}$ is solved, the optimal update is $u_{t}^{\star}=-(C+\nabla P_{R}/\rho)$, and the instantaneous Euler balance $\rho(\partial_{t}u+(u\!\cdot\!\nabla)u)=-\nabla(P_{F}+P_{R})$ is recovered.